3. Find Thevenin equivalent circuit 60 WT 0 4 AO
Answer #1

O: 61 48 Thevenin Equivalent <i> RAH → Switch OFF all independent le current source will be open ritual sources and Voltage s> s are in series. 6 and 6 Rai = 6+6 = 122 1228R 4467 RH PH Now Rri and 4 2 is in parallel RiH = Rallye = 12X4 10+4 .. RallReApply KCL at Node I 4A = If+IB » IB is current flowing through us and 62 (R2) resister as both are in So same current will fl8. Ig 6+4 15 = To = 1.5A Now UTH is Voltage across 42 Resistor = IBх 4 = 1.584 [WTH = 6 So We have RTH = 38, UTH=6V

If you have any doubt ,please feel free to ask in comment section.

Please Rate my work.